# Burnside type results for fusion categories

### v10

@@ 7. On the universal grading of hypergroups
We remark that the universal grading construction for fusion rings from [GN08, §2] works word for word in the ARN-hypergroup settings. For the completeness of the paper, we sketch its main steps below.
7.0.1. Based modules are completely reducible
@@
Given an ARN-hypergroup ${(H,\mathcal{B})}$, we define a based ${(H,\mathcal{B})}$-module as a pair $(M,\mathcal{M})$ where $M$ is a $H$-module with $\mathcal{M}$ a finite free $H$-basis for $M$, and if $x_{i}m_{j}=\sum_{k}t^{k}_{ij}m_{k}$, then $t^{k}_{ij}=t^{k^{*}}_{i^{*}j}$, for all $m_{k}\in\mathcal{M}$ and all $x_{i}\in\mathcal{B}$.
Note that the above condition is equivalent to $(L_{x_{i}^{*}})^{t}=L_{x_{i}}$, where $L_{x_{i}}$ is the matrix of the left multiplication operator by $x_{i}$ on $M$.
Then one can define a inner product on $M$ such that for all $m_{i},m_{j}\in\mathcal{M}$
(7.1) $(m_{i},m_{j})=\delta_{i,j}$
This inner product can be used to prove that any based ${(H,\mathcal{B})}$-modules is completely reducible, as in [Ost03a, Lemma 2.1].
@@
Let ${(H,\mathcal{B})}$ be a symmetric RN-hypergroup, $G$ be a finite group. Suppose that $\mathcal{B}=\sqcup_{g\in G}{\mathcal{B}}_{g}$ is a partition of $\mathcal{B}$, and let $H_{g}:=\mathbb{C}[\mathcal{B}_{g}]$. The decomposition ${(H,\mathcal{B})}=\bigoplus_{g\in G}(H_{g},\mathcal{B}_{g})$ is called a grading of ${(H,\mathcal{B})}$ by $G$ if $H_{g}H_{h}\subseteq H_{gh}$, for all $g,h\in G$. The grading is called faithful if $\mathcal{B}_{g}\neq\emptyset$, for all $g\in G$.
Let ${(H,\mathcal{B})}=\bigoplus_{g\in G}(H_{g},\mathcal{B}_{g})$ be a grading of ${(H,\mathcal{B})}$, and define $R_{g}:=\sum_{x_{i}\in\mathcal{B}_{g}}h_{i}d_{i}x_{i}$, for all $g\in G$. The analogue of [ENO05, Proposition 8.20] holds in the settings of ARN-hypergroup:
@@
Indeed, consider $R:=\sum_{g\in G}R_{g}$, and note that $\frac{1}{\mathrm{FPdim}{(H,\mathcal{B})}}R=F_{1}$, the idempotent corresponding to $\mathrm{FPdim}$ in Corollary 2.8. Thus, $RR_{h}=\mathrm{FPdim}(R_{h})R$. Since $R_{g}R_{h}\subseteq R_{gh}$, one obtains from the previous equality that $R_{g}R_{h}=\mathrm{FPdim}(R_{h})R_{gh}$, for all $g,h\in G$. Passing to $\mathrm{FPdim}$, one get that $\mathrm{FPdim}(R_{gh})=\mathrm{FPdim}(R_{g})$, for all $g,h\in G$.
@@ 7.1. Constructing the universal grading
Let ${(H,\mathcal{B})}$ be a RN-hypergroup. As above, define ${I(1)}:=\sum\limits_{a\in\mathcal{B}}h_{a}aa^{*}$. With the above notations:
@@
    A based $H_{ad}$-subbimodule $M$ of $H$ is indecomposable if and only if it is indecomposable as $H_{ad}$-bimodule.
@@
    The proof of this fact is the same as in [GN08].
@@
    Since $m{I(1)}^{n}={I(1)}^{n}m$, for all $m\in M$, this shows that both the left and the right $H_{ad}$-submodules have the same standard constituents.
@@
Let $H=\oplus_{a\in A}H_{a}$ be a decomposition of $H$ as a $H_{ad}$-bimodule. This also corresponds to a partition $\mathcal{B}=\sqcup_{a\in A}\mathcal{B}_{a}$. One may suppose that there is an element $1\in A$ such that $H_{1}=H_{ad}$.
@@ 7.1.1. Definition of $a^{*}$ for any $a\in A$
Note that if $H_{a}$ is an indecomposable $H_{ad}$-bimodule then, $(H_{a})^{*}$ is also an indecomposable $H_{ad}$-bimodule. Therefore, we may suppose that $(H_{a})^{*}=H_{a^{*}}$, for some $a^{*}\in A$. Next lemma has the same proof as [GN08, Lemma 3.4].
@@
Note that
$M:=\{y_{a}\in{\mathcal{B}}_{a}\;|\;m(y_{a},x_{a}{I(1)}^{n})>0,\;\;\text{for some}\;\;n\geq 0\}$
is a $H_{ad}$-submodule of $H_{a}$. Since $H_{a}$ is indecomposable, it follows that $M=H_{a}$. It follows that $m(y_{a}^{*},x_{a}{I(1)}^{n})>0$, which implies that $x_{a}y_{a}^{*}$ has a constituent in $H_{1}$. This gives that $x_{a}y_{a}^{*}$ is contained in $x_{a}x_{a}^{*}{I(1)}^{n}\in R_{1}.$
@@
There is a well-defined binary operation on $A$ such that $a\star b=c$ if and only if $x_{a}\star y_{b}\in H_{c}$, for all $(x_{a},y_{b})\in\mathcal{B}_{a}\times{\mathcal{B}}_{b}$. It defines a group structure on $A$, with $a^{*}$ the inverse of $a$, and $1$ the unit of $A$.
@@
Every ARN-hypergroup ${(H,\mathcal{B})}$ has a canonical faithful grading by the group $U(H)$. Any other faithful grading of ${(H,\mathcal{B})}$ by a group $G$ is determined by a surjective homomorphism $\pi:U(H)\rightarrow G$.
@@
Let ${(H,\mathcal{B})}=\oplus_{g\in G}(S_{g},\mathcal{D}_{g})$ be another grading. It follows that $xx^{*}\in\mathcal{D}_{1}$, for any $x\in\mathcal{D}_{g}$, and therefore, $\mathcal{B}_{ad}\subseteq\mathcal{D}_{1}$. Moreover, each $S_{g}$ is a ${H_{ad}}$-subbimodule of $H$. Since $H_{a}$ is an indecomposable module, then $H_{a}\subseteq S_{\pi(a)}$, for some well-defined $\pi(a)\in G$. It follows that $\pi:U(H)\rightarrow G$ is an epimorphism of groups, since we started with a grading. ∎
@@ 7.2. On the dual universal grading group and invertible
In this subsection, we use the notation $H$ for a hypergroup ${(H,\mathcal{B})}$, and $\widehat{H}$ for its dual ${\widehat{(H,\mathcal{B})}}$. The goal of this subsection is to prove the following:
@@
Let $H$ be an abelian RN-hypergroup. Then,
$U_{H}\simeq{G(\widehat{H})}$
Dually one has the following:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then
$U_{\widehat{H}}\simeq G(H)$
as abelian groups.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying any of the two hypothesis from Lemma 2.23. Recall that this means that al least one of the following holds:
  1. (1)
    ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$
  2. (2)
    ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
By Corollary 2.32, for any such hypergroup, if $\mu_{j}\in{G(\widehat{H})}$ then $\mu_{j}^{n}=\mu_{1}$, for some $n\geq 0$. Thus, for all $i$, $\mu_{j}(x_{i})=\omega_{ij}d_{i}$ for some complex number $|\omega_{ij}|=1$.
@@
Let ${(H,\mathcal{B})}$ be a hypergroup satisfying any of the two hypothesis from Lemma 2.23. For any standard element $x_{i}\in\mathcal{B}$, we define a morphism $\omega_{i}:{G(\widehat{H})}\rightarrow\mathbb{C},\mu_{j}\mapsto\omega_{ij}=\mu_{j}({\frac{x_{i}}{d_{i}}})$.
Next lemma shows that $\omega_{i}$ is a character of ${G(\widehat{H})}$.
Lemma 7.11.
Let ${(H,\mathcal{B})}$ be a hypergroup satisfying any of the two hypothesis from Lemma 2.23. For any $x_{i}$, then $\omega_{i}$ is a linear character of ${G(\widehat{H})}$.
@@
Since ${\frac{x_{i}}{d_{i}}}:\widehat{H}\rightarrow\mathbb{C}$ are algebra characters, then their restrictions ${\frac{x_{i}}{d_{i}}}\big{|}_{{G(\widehat{H})}}$ are also algebra characters. Note that ${\frac{x_{i}}{d_{i}}}\big{|}_{{G(\widehat{H})}}=\omega_{i}$. ∎
Let ${(H,\mathcal{B})}$ be a satisfying any of the two hypothesis from Lemma 2.23. Given $\psi\in\widehat{{G(\widehat{H})}}$, let
(7.3) $\mathcal{B}_{\psi}:=\{x_{i}\in\mathcal{B}\;\mid\;\omega_{i}=\psi\}=\{x_{i}\;|\;\mu_{j}({\frac{x_{i}}{d_{i}}})=\psi(\mu_{j}),\;\text{for all}\;\mu_{j}\in{G(\widehat{H})}\}$
Note that the above definition can be rewritten as
$x_{i}\in\mathcal{B}_{\psi}\iff{\frac{x_{i}}{d_{i}}}\big{|}_{G(\widehat{H})}=\psi$
Proposition 7.12.
Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup. Then, Equation (7.3) defines a grading on ${(H,\mathcal{B})}$.
Proof.
It reduces to show that if $x_{i}\in\mathcal{B}_{\psi}$ and $x_{k}\in\mathcal{B}_{\phi}$, then $x_{i}x_{k}$ has all constituents in $\mathcal{B}_{\psi\phi}$. Note that $\mu_{j}({\frac{x_{i}}{d_{i}}})=\psi(\mu_{j})$ and $\mu_{j}(\frac{x_{k}}{d_{k}})=\phi(\mu_{j})$, for all $\mu_{j}\in{G(\widehat{H})}$. It follows that
@@
Now, since $|\psi\phi(\mu_{j})|=1$ and ${\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}}=\frac{1}{d_{i}d_{k}}\sum_{l=1}^{m}d_{l}N^{l}_{ik}\frac{x_{l}}{d_{l}}$, the triangle inequality implies that $\mu_{j}(\frac{x_{l}}{d_{l}})=\mu_{j}(\frac{x_{i}}{d_{i}})\mu_{j}(\frac{x_{k}}{d_{k}})$, for all $l$ with $N^{l}_{ik}>0$. Indeed, note that
$1=|\psi\phi(\mu_{j})|=|\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}})|\leq\frac{1}{d_{i}d_{k}}\sum_{l=1}^{m}d_{l}N^{l}_{ik}|\mu_{j}(\frac{x_{l}}{d_{l}})|\leq\frac{1}{d_{i}d_{k}}\sum_{l=1}^{m}d_{l}N^{l}_{ik}=1,$
and this forces $\mu_{j}(\frac{x_{l}}{d_{l}})=\mu_{j}(\frac{x_{i}}{d_{i}})\mu_{j}(\frac{x_{k}}{d_{k}})$. ∎
7.3. Computations of the regular components
For any grading component $\mathcal{B}_{\psi}$, we define (as above) the regular part of the grading:

### v11

@@ 7. On the universal grading of hypergroups
The universal grading construction for fusion rings from [GN08, §2] works word for word in the ARN-hypergroup settings. For the completeness of the paper, we sketch its main steps below.
7.1. Based modules are completely reducible
@@
Given an ARN-hypergroup ${(H,\mathcal{B})}$, we define a based left ${(H,\mathcal{B})}$-module as a pair $(M,\mathcal{M})$ where $M$ is a $H$-module with $\mathcal{M}$ a finite free $H$-basis for $M$, i.e. for all $x_{i}\in\mathcal{B}$ and all $m_{j}\in\mathcal{M}$, then $x_{i}m_{j}=\sum_{k}t^{k}_{ij}m_{k}$; and moreover, for all $i,j,k$ then $t^{k}_{ij}$ is zero if and only if $t^{j}_{i^{*}k}$ is zero.
The last condition above means that $(x_{i}m_{j},m_{k})$ vanishes if and only if $(m_{j},x_{i^{*}}m_{k})$ vanishes, for the inner product on $M$ such that for all $m_{i},m_{j}\in\mathcal{M}$
(7.1) $(m_{i},m_{j})=\delta_{i,j}.$
It is easy to deduce that any based ${(H,\mathcal{B})}$-modules is completely reducible, as in [Ost03a, Lemma 2.1]. Indeed, if $N\subset M$ and $HN\subset N$ then $(N,N^{\perp})=\{0\}=(HN,N^{\perp})=(N,HN^{\perp})$. Thus, $HN^{\perp}\subset N^{\perp}$.
@@
Let ${(H,\mathcal{B})}$ be a RN-hypergroup, $G$ be a finite group. Suppose that $\mathcal{B}=\sqcup_{g\in G}{\mathcal{B}}_{g}$ is a partition of $\mathcal{B}$, and let $H_{g}:=\mathbb{C}[\mathcal{B}_{g}]$. The decomposition ${(H,\mathcal{B})}=\bigoplus_{g\in G}(H_{g},\mathcal{B}_{g})$ is called a grading of ${(H,\mathcal{B})}$ by $G$ if $H_{g}H_{h}\subseteq H_{gh}$, for all $g,h\in G$. The grading is called faithful if $\mathcal{B}_{g}\neq\emptyset$, for all $g\in G$.
Please note that a faithful grading group of an abelian hypergroup must also be abelian. Let ${(H,\mathcal{B})}=\bigoplus_{g\in G}(H_{g},\mathcal{B}_{g})$ be a faithful grading of ${(H,\mathcal{B})}$, and define $R_{g}:=\sum_{x_{i}\in\mathcal{B}_{g}}h_{i}d_{i}x_{i}$, for all $g\in G$. The analogue of [ENO05, Proposition 8.20] holds in the settings of ARN-hypergroup:
@@
Indeed, consider $R:=\sum_{g\in G}R_{g}$, and note that $\frac{1}{\mathrm{FPdim}{(H,\mathcal{B})}}R=F_{1}$, the idempotent corresponding to $\mathrm{FPdim}$ in Corollary 2.10. Thus, $RR_{h}=\mathrm{FPdim}(R_{h})R$. Since $H_{g}H_{h}\subseteq H_{gh}$, we obtain from the previous equality that $R_{g}R_{h}=\mathrm{FPdim}(R_{h})R_{gh}$, for all $g,h\in G$. Passing to $\mathrm{FPdim}$, we get that $\mathrm{FPdim}(R_{g})\mathrm{FPdim}(R_{h})=\mathrm{FPdim}(R_{h})\mathrm{FPdim}(R_{gh})$; but $\mathrm{FPdim}(R_{h})\neq 0$ by faithfulness, so $\mathrm{FPdim}(R_{gh})=\mathrm{FPdim}(R_{g})$, for all $g,h\in G$.
@@ 7.2. Constructing the universal grading
Let ${(H,\mathcal{B})}$ be a RN-hypergroup. For all $x\in H$, define $I(x):=\sum_{j\in\mathcal{I}}h_{j}x_{j}xx_{j^{*}}$. With the above notations:
@@
    A based $H_{ad}$-subbimodule $M$ of $H$ is indecomposable if and only if it is indecomposable as left $H_{ad}$-module.
@@
    For all $i\in\mathcal{I}$, by Frobenius reciprocity (2.18):
    $x_{i}I(x)=\sum_{j\in\mathcal{I}}h_{j}x_{i}x_{j}xx_{j^{*}}=\sum_{j,k\in\mathcal{I}}h_{j}N_{i,j}^{k}x_{k}xx_{j^{*}}=\sum_{j,k\in\mathcal{I}}h_{k}N_{k^{*},i}^{j^{*}}x_{k}xx_{j^{*}}$
    $=\sum_{k\in\mathcal{I}}h_{k}x_{k}x(\sum_{j\in\mathcal{I}}N_{k^{*},i}^{j^{*}}x_{j^{*}})=\sum_{k\in\mathcal{I}}h_{k}x_{k}xx_{k^{*}}x_{i}=I(x)x_{i}.$
@@
    Recall that $H_{ad}:=\langle I(1)\rangle$. By (1), $m{I(1)}^{n}={I(1)}^{n}m$, thus by RN-assumption, $mH_{ad}=H_{ad}m\subset M$, for all $m\in M$, meaning that $M$ is also a based right $H_{ad}$-submodule.
@@
Let $H=\oplus_{a\in A}H_{a}$ be a decomposition (into indecomposable) of $H$ as a $H_{ad}$-bimodule. This also corresponds to a partition $\mathcal{B}=\sqcup_{a\in A}\mathcal{B}_{a}$. Let 111 be an element of $A$ such that $H_{1}=H_{ad}$.
@@ 7.2.1. Definition of $a^{*}$ for any $a\in A$
Since $H_{a}$ is an indecomposable $H_{ad}$-bimodule then $(H_{a})^{*}$ is also an indecomposable $H_{ad}$-bimodule. Let $a^{*}$ be an element of $A$ such that $(H_{a})^{*}=H_{a^{*}}$. Next lemma is as [GN08, Lemma 3.4].
@@
Note that $M:=H_{a}\cap(H_{ad}y_{a})$ is a $H_{ad}$-submodule of $H_{a}$ (containing $y_{a}$). But $H_{a}$ is indecomposable (so irreducible by complete reducibility). It follows that $M=H_{a}$, so $H_{a}\subset H_{ad}y_{a}$. Thus $x_{a}\in H_{ad}y_{a}$ and $x_{a}y_{a}^{*}\in H_{ad}y_{a}y_{a}^{*}\subset H_{ad}$. ∎
@@
For all $a,b\in A$, there is $c\in A$ such that $H_{a}H_{b}\subset H_{c}$. This induces a group structure on $A$ given by $ab:=c$. Moreover, $a^{*}$ is the inverse of $a$, and 111 the unit of $A$.
@@
Every ARN-hypergroup ${(H,\mathcal{B})}$ has a canonical faithful grading by the group $U(H)$. Any other faithful grading of ${(H,\mathcal{B})}$ by a group $G$ is determined by a surjective group morphism $\pi:U_{H}\rightarrow G$.
@@
Let ${(H,\mathcal{B})}=\oplus_{g\in G}(S_{g},\mathcal{D}_{g})$ be another faithful grading. It follows that $xx^{*}\in S_{1}$, for any $x\in\mathcal{D}_{g}$, but $xx^{*}\in H_{ad}$ therefore, $\mathcal{B}_{ad}\subseteq\mathcal{D}_{1}$. Moreover, each $S_{g}$ is a ${H_{ad}}$-subbimodule of $H$. Since $H_{a}$ is an indecomposable module, then $H_{a}\subseteq S_{\pi(a)}$, for some well-defined $\pi(a)\in G$. It follows that $\pi:U_{H}\rightarrow G$ is a surjective group morphism. Indeed, for all $a,b\in A$, then $H_{a}\subset S_{\pi(a)}$ and $H_{b}\subset S_{\pi(b)}$, thus $H_{a}H_{b}\subset S_{\pi(a)}S_{\pi(b)}\subset S_{\pi(a)\pi(b)}$. But $H_{a}H_{b}\subset H_{ab}$, thus $H_{a}H_{b}\subset H_{ab}\cap S_{\pi(a)\pi(b)}\neq\emptyset$, so $H_{ab}\subset S_{\pi(a)\pi(b)}$, meaning that $\pi(ab)=\pi(a)\pi(b)$. Finally, the surjectivity comes from the faithfulness. ∎
@@ 7.3. On the dual universal grading group and invertible
In this subsection, we use the notation $H$ for a hypergroup ${(H,\mathcal{B})}$, and $\widehat{H}$ for its dual $^{\widehat{(H,\mathcal{B})}}$. The goal of this subsection is to prepare for the proof (§7.4) of the following:
@@
Let $H$ be an ARN-hypergroup. Then,
$U(H)\simeq{G(\widehat{H})}$
Dually, we have the following:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then
$U(\widehat{H})\simeq G(H).$
Recall from Lemma 2.28 that for any abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ then $\mu_{j}\in{G(\widehat{H})}$ if and only if $|\omega_{ij}|=1$, for all $i\in\mathcal{I}$, with $\omega_{ij}:=\mu_{j}(\frac{x_{i}}{d_{i}})$.
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. For any $x_{i}\in\mathcal{B}$, we define the linear character $\omega_{i}:={\frac{x_{i}}{d_{i}}}\big{|}_{{G(\widehat{H})}}$, the restriction to ${G(\widehat{H})}$ of the linear character $^{\frac{x_{i}}{d_{i}}}:\widehat{H}\rightarrow\mathbb{C}$, i.e. $\omega_{i}(\mu_{j})=\omega_{ij}$.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Given $\psi\in\widehat{{G(\widehat{H})}}$, let
(7.3) $\mathcal{B}_{\psi}:=\{x_{i}\in\mathcal{B}\;\mid\;\omega_{i}=\psi\}=\{x_{i}\;|\;\mu_{j}({\frac{x_{i}}{d_{i}}})=\psi(\mu_{j}),\;\text{for all}\;\mu_{j}\in{G(\widehat{H})}\}.$
Proposition 7.11.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Then $(\mathcal{B}_{\psi})_{\psi\in\widehat{{G(\widehat{H})}}}$ defines a grading on ${(H,\mathcal{B})}$.
@@
For any $\mu_{j}\in{G(\widehat{H})}$, $\psi,\phi\in\widehat{{G(\widehat{H})}}$, $x_{i}\in B_{\psi}$ and $x_{k}\in B_{\phi}$, then
@@
By the triangle inequality:
$1=|\psi\phi(\mu_{j})|=|\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}})|\leq\frac{1}{d_{i}d_{k}}\sum_{l\in\mathcal{I}}d_{l}N^{l}_{ik}|\mu_{j}(\frac{x_{l}}{d_{l}})|\leq\frac{1}{d_{i}d_{k}}\sum_{l\in\mathcal{I}}d_{l}N^{l}_{ik}=1,$
which forces $\mu_{j}(\frac{x_{l}}{d_{l}})=\mu_{j}(\frac{x_{i}}{d_{i}})\mu_{j}(\frac{x_{k}}{d_{k}})$ if $N^{l}_{ik}$ is nonzero, meaning that if $x_{i}\in\mathcal{B}_{\psi}$ and $x_{k}\in\mathcal{B}_{\phi}$, then $x_{i}x_{k}$ has all constituents in $\mathcal{B}_{\psi\phi}$. ∎
7.4. Computation of the regular components
For any grading component $\mathcal{B}_{\psi}$, we define (as above):